OpenStudy (anonymous):

8.38888889 as a fraction plz thx

OpenStudy (anonymous):

838888889/10^8

OpenStudy (anonymous):

8.38888889 that's rounded form of \(8.38888888888\cdots \) isn't ?

myininaya (myininaya):

x=8.3888888888888888888888888888 10x=83.888888888888888888888888888888 10x-x=83.88-8.38 9x=75.5 90x=755 x=755/90

myininaya (myininaya):

yes you could

myininaya (myininaya):

if you notice the top number ends in a 5 and the bottom numbers end in a 0 so both numbers are divisible by 5

OpenStudy (anonymous):

151/18

OpenStudy (anonymous):

o ok thx i see

myininaya (myininaya):

and then it can be reduced more as fontez as shown

OpenStudy (anonymous):

151/18

OpenStudy (anonymous):

\[ 8 \frac{35}{90} \]

myininaya (myininaya):

or written as an improper fraction as ffm has shown

OpenStudy (anonymous):

I used a quick trick but if you want to understand the analytical/ formal way of doing it see myin's answer.